### v1

# Toward Agentic Software Engineering Beyond Code: Framing Vision, Values, and Vocabulary

Abstract.
Agentic AI is poised to usher in a seismic paradigm shift in Software Engineering (SE). As technologists rush head-along to make agentic AI a reality, SE researchers are driven to establish agentic SE as a research area. While early visions of agentic SE are primarily focused on code-related activities, early empirical evidence calls for a consideration of a range of socio-technical concerns to make it work in practice. This paper contributes to the emerging community vision by: (a) recommending an expansion of its scope beyond code, toward a ‘whole of process’ vision, grounding it in SE foundations and evolution and emerging agentic SE frameworks, (b) proposing a preliminary set of values and principles to guide efforts, and (c) sharing guidance on designing/using well-defined vocabulary for agentic SE. It is hoped that these ideas will encourage community collaborations and steer the SE community towards laying strong foundations of agentic SE so its not only inevitable but also deliberate and desirable in the long run.
Agentic software engineering, process, vision, values, principles, vocabulary, terminology, Agentic AI
This version is the authors’ preprint, submitted to arXiv on 22-Oct-2025.
1. Introduction
The Software Engineering (SE) industry, and consequently the SE research discipline, is at the crux of an incoming seismic paradigm shift proposed by the highly anticipated advent of agentic AI. Latest releases of agentic systems (e.g., SWE-agent, Passerine, MASAI AutoDev, Devin) are fast making agentic AI a reality. Following suit, visions capturing various components of agentic SE are being proposed, e.g., agentic AI software engineers (Roychoudhury et al., 2025), unified software engineering agent (USEagent) (Applis et al., 2025), and structured agentic software engineering (SASE) (Hassan et al., 2025). Li et al. have made AIDev dataset comprising of over 450K agentic pull requests created by leading autonomous coding agents publicly available (Li et al., 2025), positioning agentic SE as an “unfolding reality”. These visions, while necessary and welcome, are primarily focused on one SE activity – programming.
If we expand our vision to focus on SE activities beyond the necessary technical foundations, a range of equally critical socio-technical concerns emerge. Early empirical studies show AI’s role as a “personal accelerator’ for coding, writing, and documentation tasks but highlight its limitations in addressing teamwork, coordination, accountability, and culture (Xiao and et al., 2025). Similarly, cross-phase applicability, organizational adoption challenges, and human-AI collaboration have been highlighted as areas deserving research focus (Akbar et al., 2025). In a review of AI agentic programming, Wang et al. highlight the increasing human and socio-technical concerns requiring interdisciplinary solutions (Wang et al., 2025). Abrahão et al. pose several socio-technical questions for the age of AI (e.g., role of human software engineers, how human-AI collaboration might work, how “citizen engineers” might interact with AI agents) (Abrahão et al., 2025).
There is an urgent need to expand our emerging vision of agentic SE beyond code to examine a full range of SE activities, artefacts, roles, and concerns – traditional ones renewed for agents and multi-agent systems and new ones required to make human-agent ecosystems work. This paper contributes to the emerging agentic SE vision by: (a) extending its scope beyond coding, toward a ‘whole of process’ vision, by grounding it in a discussion of SE foundations and evolution (in Section 2), (b) proposing a preliminary set of values and principles to guide our collective design of the emerging vision (in Section 4), and (c) highlighting the importance of, and sharing guidance on, building well-defined vocabulary (terminology) for the emerging agentic SE area (in Section 5). Together, it is hoped that these ideas will steer our emerging community vision of agentic SE in the direction of laying the core foundation of agentic SE process models, more fleshed out values and principles, and well-defined vocabulary to support our scholarly conversations, in verbal and written forms. As with other recent visions, these are not meant to be exhaustive or final. Rather, they are intended to encourage community feedback and discussions about adopting a systematic and principled approach to thinking about, researching, and enabling agentic SE.
2. SE Foundations and Evolution
Before we can redefine a field, it makes sense to revisit its foundational definition and evolution. SE is defined as “the application of a systematic, disciplined, quantifiable approach to the development, operation, and maintenance of software; that is, the application of engineering to software.” (ISO/IEC/IEEE 24765:2017 Systems and Software Engineering — Vocabulary/Guide to SWEBOK 4.0a [^1] https://ieeecs-media.computer.org/media/education/swebok/swebok-v4.pdf)
Ergo, SE is more than programming. Starting out as a “cottage industry” in the 1950s, SE became an established engineering discipline with many SE process models (AKA software development lifecycle (SDLC), frameworks, methods) proposed, adopted, established, and replaced over the decades (Hoda et al., 2018). All of them featured a ‘whole of process’ approach. Royce introduced structure and formality to SE in 1970 through a two iteration approach called Waterfall. It was widely adopted as a specification-driven, sequential approach in practice, featuring core SE activities as phases (requirements, analysis, design, coding, testing, and operations), artefacts (documents from each phase), and formalized roles (requirements engineers, designers, developers, testers etc.). Over time, tools and techniques were introduced and refined to support the roles to implement the activities and produce the artefacts. Later, the Sprial model and V-model espoused a risk-driven and verification-validation-driven approach to SE, respectively. While, the Rational Unified Process (RUP) highlighted software project management concerns such as timely delivery and budget (Kruchten, 2004).
Agile software development marked a paradigm shift in SE with its people-centeredness, collaborative approach, focus on working software, and ability to pivot in response to changes (Manifesto, 2001). Agile methods, eXtreme Programming and Scrum, came to dominate the SE industry (Hoda et al., 2018). Agile SE researchers studied its newly introduced roles, practices, artefacts, its impact on related (sub)disciplines (e.g., requirements engineering, project management) and integration with seemingly contradictory approaches (e.g., CMMI, safety-critical systems) (Hoda et al., 2018). Lean principles manifested through Kanban. Extending the coverage of the SDLC, DevOps connected development and operations. Agile became a default/standard (Digital.ai, 2024).
With the recent resurgence of AI, there was renewed interest in augmenting SE with AI support. Dam et al. (Dam et al., 2019), proposed a conceptual framework of AI/ML powered agile. Building on that, Hoda et al. proposed Augmented Agile, an AI-powered agile project management approach, combining the ‘human heart’ (e.g., humans values and concerns) with ‘AI brains’, imagined as an agile co-pilot acting as a team member (Hoda et al., 2023). However, these efforts remain conceptual, without application in industrial contexts. They foreshadowed what became GenAI/LLM powered SE practices, such as LLM-based multi-agent systems for SE (He et al., 2025). In a review of 395 papers focused on the use of LLM-based solutions for SE challenges, Hou et al. found a strong focus on individual SE activities, with software development, (programming or coding), being the most researched SE activity (Hou et al., 2024). Recently, Zhang et al. proposed AgileGen, an agile-based generative software development through human-AI teamwork framework, focused on maintaining consistency between the generated software and user requirements (Zhang and et al., 2025). While others have called for GenAI to be human-centered (Russo et al., 2024).
Figure 1. Software Engineering Evolution: A Brief Timeline
There is an urgent need to expand our vision of agentic SE beyond code (design, writing, testing, reviewing, repairing) to examine a full range of agentic SE areas and associated actors, activities, artefacts, and other socio-technical concerns – traditional SE ones renewed for agents and multi-agent systems and new ones needed to make human-agent ecosystems work. The establishment of SE can be credited to the formalization of various process models, the development of supporting tools, techniques, and the formalization of knowledge areas and vocabulary. We need the same to establish agentic SE.
3. Toward Agentic SE (emerging)
Agentic AI draws its foundations from agent-oriented software engineering (Wooldridgey and Ciancarini, 2000; Hassan et al., 2025; Roychoudhury, 2025). Introduction of agentic systems such as SWE-agent, Google’s Passerine, Microsoft’s MASAI, AutoDev, Devin are making agentic AI a reality. Unsurprisingly, visions of new agentic SE frameworks and approaches are being proposed.
In an opinion piece, Roychoudhury et al. propose a new role of agentic AI Software Engineer, an autonomous LLM-based coding agent “that combines coding, testing, debugging, etc. into a coherent, explainable workflow(Roychoudhury et al., 2025). They predict a shift in interest from programming in the large to programming with trust and highlight the importance of operationalizing socio-technical integration of AI agent trust, predicting a future where human developers can rely on autonomous agents to become trusted programming collaborators. The idea of a unified software engineering agent (USEagent) with capabilities to handle multiple programming tasks (e.g., CodeRetrieval, ExecuteTests, EditCode, ReviewPatch) is described in (Applis et al., 2025). It promises to be “a first draft of a future AI Software Engineer which can be a team member in future software development teams involving both AI and humans.
Hassan et al. propose structured agentic software engineering (SASE) as a conceptual framework for “two symbiotic modalities”, namely, SE for Humans and SE for Agents (Hassan et al., 2025). They propose a hierarchical framework with increasing levels of autonomy: from manual coding (level 0) to goal-agentic (dubbed Agentic SE/SE 3.0) and all the way to general domain autonomy (level 5, SE 5.0). They envision humans and agents engaging in agent-based software programming activities supported by dedicated agent command environment (ACE) and agent execution environment (AEE). They share a vision of agents driving software programming, supported by evolving artefacts (e.g., briefingpacks), roles (e.g., humans as coaches and mentors for agents) and tools (e.g., ACE and AEE replacing traditional IDEs). Recently, Li et al. made AIDev, a large-scale dataset of over 450K PRs by leading autonomous agents, publicly available for SE and AI researchers to longitudinally study a range of research questions, e.g., to do with agent behaviour and dynamics, human-AI collaboration, productivity, effort, quality etc. (Li et al., 2025). These visions and efforts, while necessary and welcome, are primarily focused on one SE activity – programming.
Besides the technical challenges of implementing agentic AI for programming-related SE tasks, a number of socio-technical concerns across the SDLC arise. Early empirical evidence highlight the need to address agentic AI applicability across the SE process. Xiao et al.’s longitudinal study showed AI’s role as ‘a personal accelerator’ but it was not seen to fix teamwork issues and its impact on coordination, accountability, and culture remain to be seen (Xiao and et al., 2025). Akbar et al. report that current studies “often focus narrowly on AI-powered code generation or bug detection without addressing”: cross-phase applicability, organizational adoption challenges, and human-AI collaboration patterns (Akbar et al., 2025). Similarly, Wang et al. review AI agentic programming and highlight that increasing human and socio-technical concerns with AI agentic programming such as context handling, human-AI collaboration, verification of agent behaviour, will require “interdisciplinary solutions(Wang et al., 2025). Similar socio-technical issues (e.g., human-AI collaboration, next generation education, job security, ethics, diversity) were highlighted by Abrahão et al. (Abrahão et al., 2025). How nascent agentic SE is can be gauged from the fact that most of these proposals are under review (Oct 2025).
A ‘Whole of Process’ Vision
Building on the foundational aspects of previous SE models (Section 2) and AI capabilities captured in proposed frameworks above, this paper presents a preliminary ‘whole of process’ vision for agentic SE (Figure 2). It is not meant to be a definitive agentic SE process model, rather a ‘first look’ that will need to be improved and expanded on. The vision extends across high-level core areas, approached iteratively, by human and agent (single AI agents or a multi-agent system) actors, collaborating in activities at varying levels of AI agency (depicted by the slider under human control).
  • Agentic SE Ethical Alignment: covers activities, actors, and artefacts (AAA) to do with AI and software ethics alignment. With the increasing capabilities and agency of AI models and their potential for wide-scale impact, ethical alignment is a must.
  • Agentic Requirements Engineering: covers AAA to do with software requirements (customer, user, subject matter expert, quality, and ethical requirements)
  • Agentic Design: covers AAA to do with software design (architecture, UI, UX).
  • Agentic Development: covers AAA to do with software code design, production, testing, reviewing, repairing. We have a head-start on envisioning the transformation in this area with emerging visions and an agentic dataset. (Roychoudhury et al., 2025; Roychoudhury, 2025; Hassan et al., 2025; Li et al., 2025).
  • Agentic Operations: covers AAA to do with deployment, maintenance, and operations (e.g., Ops/DevOps/DevSecOps).
These high-level descriptions above need inputs from relevant area experts to develop details of how the vision may be implemented in practice. Transforming relevant SE areas and introducing new ones (e.g., ethical alignment) will ensure agentic SE leverages the full capabilities of agents to support humans and improve outcomes across the process, not only to do with code-related activities.
Figure 2. A ‘whole of process’ vision for Agentic SE, where human and agent actors iterate across areas, activities, artefacts with varying levels of AI agency (under human control)
4. Toward Agentic SE Values and Principles
To guide this paradigm shift, a set of preliminary CRAFT values and principles for agentic SE is presented to spark community conversation and collaboration.
Value 1. Comprehensive:
Agentic SE needs to take a ‘whole of process approach’ and include the consideration of human and socio-technical aspects in addition to core technological innovation.
Table 1. CRAFT Values and Principles for Agentic Software Engineering (Preliminary)
Value Principle For Agentic SE
Comprehensive Whole of Process Approach Agentic SE needs to take a ‘whole of process’ approach to examining traditional SE areas and associated activities, actors, artefacts (AAA) and consider new ones to enable emerging human-agent ecosystems.
Human and Socio-technical Aspects Agentic SE needs to include a focus on human and socio-technical aspects of SE including roles played by human and agent actors, human aspects that apply to humans and agents (e.g., human values, empathy, motivation, emotional intelligence, personality) and socio-technical aspects (e.g., trust, adoption barriers, teamwork, collaboration, communication, coordination).
Responsible Ethics-by-Design Agentic SE needs to needs to apply an ethics-by-design approach to its design and application, questioning the why, and not being driven by the why not.
Sustainability Agentic SE need to ensure sustainability of the core technologies, ecosystems, and the planet.
Adaptive Industrial Relevance Agentic SE need to adapt to new AI models and paradigms and work closely with the AI community and co-design with the industry.
Socio-technical Relevance Agentic SE needs to predict and report socio-technical impacts of changes in AI models and paradigms.
Foundational Foundational Knowledge Agentic SE need to develop foundational knowledge about AAA across all areas, not only programming.
Foundational Solutions Agentic SE need to develop foundational solutions to support AAA across all areas, not only programming.
Translational Awareness and Education Agentic SE need to build awareness through education, certifications, funding bodies, media, and society; and maintain incident logs.
Actionable Guidelines Agentic SE need to study and enable the adoption of research knowledge and solutions into practice through usable guidelines and tools.
  • Principle of Whole of ProcessAgentic SE needs to take a ‘whole of process’ approach – A ‘whole of process’ approach, a hallmark of previous SE models (section 2). Perhaps being one of the reasons why Scrum (covering the full SE lifecycle and project management aspects) considerably outperformed extreme programming (Hoda et al., 2018). Relevant parts of the SE lifecycle, not only its code-related activities, need to be considered as well as new required activities, in designing agentic SE. This includes assessing the goals and constrains related to ideation, RE activities (e.g., to establish user needs and preferences), design, architecture, project management (e.g., costs, times, resources), functional and user testing, delivery, maintenance, operations/evolution, as well as to new practices, trade-offs, and process optimization.
  • Principle of Human and Socio-technical AspectsAgentic SE needs to include the human and socio-technical aspects of SE – Like SE, agentic SE is socio-technical, where “the social and technical aspects are inherently interwoven(Hoda, 2024). Technical innovation is necessary but not sufficient for advancing agentic SE, e.g., trust and utility concerns outrank performance in real-world adoption (Li et al., 2025). Human and socio-technical aspects must be considered, e.g., the roles played by human and agent actors, human aspects (e.g., human values, empathy, emotional intelligence), and socio-technical aspects (e.g., trust, adoption barriers, human experience, coordination, communication, collaboration, conflict management, teamwork). Research on these will help the community answer fundamental questions such as: What should be automated, to what extent (human oversight levels), and how? What shouldn’t be automated and why? How should agentic systems be best integrated into teams and workflows?
Value 2. Responsible:
Agentic SE needs to ensure ethics-by-design and sustainability. With agentic AI injecting more agency into AI, we need to aim for positive impact and human control of AI (Spiegler et al., 2025).
  • Principle of Ethics-by-DesignAgentic SE needs to apply an ethics-by-design approach to its design and application – There is unprecedented growth in AI-based companies (e.g., OpenAI, NVIDIA, Microsoft, Cursor) with claims of 100x productivity [^2] https://www.businessinsider.com/surge-ceo-ai-100x-engineers-2025-7. But to what end and at what cost? It is critical to take an ethics first approach to agentic SE, where ethical aspects (e.g., accountability, reliability, safety, transparency, explainability, privacy, security, human/societal/environmental wellbeing (Lu et al., 2022)) are prioritized. Applying ethical research, we should question the why and not be driven by the why not (Storey et al., 2025) (e.g., why does X need to be automated vs. why not automate X).
  • Principle of SustainabilityAgentic SE needs to ensure sustainability – The planetary cost of AI technologies is non-trivial (Crownhart, 2025). As Cruz et al. highlight in their article on Green AI, “our [SE] community needs to adopt a long-term plan enabling a conscious transformation that aligns with environmental sustainability values(Cruz et al., 2025), which requires handling trade-offs (Lago et al., 2015).
Value 3. Adaptive:
Agentic SE needs to adapt to maintain industrial and socio-technical relevance.
  • Principle of Industrial RelevanceAgentic SE researchers need to adapt to new AI models and paradigmsAgentic SE fundamentally ties SE to AI. Agentic SE and AI communities should work in tandem, adapting and contributing to the latest AI models and paradigms. Empirical studies of problem and solution spaces will be useful in maintaining industrial relevance. Co-designing and evaluations with industry will also help.
  • Principle of Socio-technical RelevanceAgentic SE needs to predict and report socio-technical impacts of changes in AI models and paradigms – Maintaining relevance extends to the associated human (e.g., evolving levels of AI agency and human control (Spiegler et al., 2025)) and socio-technical aspects (e.g., awareness, ethics, adoption). This implies anticipating and keeping up with the impacts of changing AI technology and envisioning, studying, and reporting them promptly to inform research and practice.
Value 4. Foundational:
Agentic SE needs to develop foundational knowledge and solutions.
  • Principle of Foundational KnowledgeAgentic SE needs to develop foundational knowledge – Like how agile established foundational knowledge through empirical studies of new roles, practices, artefacts, and socio-technical aspects (Hoda et al., 2018), a similar approach will benefit agentic SE. Building original taxonomies and theories (Hoda, 2024) and testing existing ones will contribute to theoretical foundations, e.g., agents in SE taxonomy (Wang and et al., 2025). Qualitative (Seaman et al., 2025), quantitative, and mixed methods studies (Storey et al., 2025) will help build a body of knowledge.
  • Principle of Foundational SolutionsAgentic SE needs to develop foundational solutions – With proposed visions of programming frameworks (Hassan et al., 2025; Roychoudhury et al., 2025), there is already movement on this front. As with GenAI/LLM based solutions (technical tools, techniques, frameworks, platforms) for SE (Hou et al., 2024), we can expect foundational solutions for agentic SE being proposed by the SE community. It will be critical to extend the solutions to support other parts of agentic SE areas beyond programming.
Value 5. Translational:
Agentic SE needs to enable research translation into practice through building awareness, education, and actionable guidelines.
  • Principle of Awareness and EducationAgentic SE researchers need to build awareness – While awareness of agentic AI is on the rise thanks to agentic AI systems, products, and news (including ethical incidents), awareness of agentic SE will need to be built through education and training within and across disciplines (in classrooms and workplaces) as well as with accreditation, certification, funding bodies, media, and society. Maintaining an agentic SE incident log will also help. Previous SE models have shown the importance of such efforts to the long-term future of those models and their associated communities.
  • Principle of Usable Guidelines and ToolsAgentic SE researchers need to study and enable the adoption of research knowledge and solutions into practice – Building knowledge and solutions is necessary but not sufficient to progress agentic SE. To achieve real-world impact, they need to be translated into usable guidelines and tools for practice. Mirroring and leveraging AI’s multi-modal generation, they can have more industry-friendly interfaces and mode of delivery (e.g., posters, executive summaries, graphical abstracts, short videos).
5. Toward Agentic SE Vocabulary
Well-defined terminology is a precursor to formalizing taxonomy (e.g. Sapkota et al. distinguish between AI agents and agentic AI and present a taxonomy of AI agent paradigms (Sapkota et al., 2025)) and ultimately, knowledge areas. While agentic SE is nascent, discrepancies in terminology usage are already apparent. Some are syntactic (e.g., agentic AI software engineer vs. AI software engineer vs. agentic software engineer). While others are reflective of deeper semantic and even philosophical issues. So we can lay solid foundations for agentic SE, here are some considerations for designing and using terminology for agentic SE:
  • Relevancefit for purpose and domain context – to what extent is the term relevant to the domain? E.g., traditional SE vocabulary can be revisited for relevance in the agentic SE context.
  • Coverageconceptual completeness and accuracy – to what extent does the term cover (nearly) all the goals and responsibilities of the agentic SE process area, actor, activity, or artefact it aims to support or automate? For example, a software engineer is more than a programmer, so can an agent that enacts programming tasks be termed a software engineer? Autonomous coding agent (Li et al., 2025) is a more accurate description.
  • Acceptancerecognition and adoption by the community – to what extent will humans (especially in the role) agree to an AI agent being given their professional title? For example, people might judge human and robot paintings as art but are less willing to see robots as artists (Mikalonytė and Kneer, 2022).
  • Consistencyuniformity of use, stability, adaptability – to what extent do different research groups around the world within SE use the terminology in the same way, both in form (syntax) and meaning (semantics)? how long do they remain stable? when and how do they adapt to changes over time?
  • Philosophical alignmentreflexivity leading to awareness of underlying assumptions, power dynamics, fairness – reflecting on who comes up with terminology and why? Who stands to gain, who stands to lose, and does everyone have a voice (e.g., global south, underfunded regions)? How can we ensure fairness?
6. A Deliberate and Desirable Paradigm Shift
Building on the foundational aspects of previous SE models (Section 2) and recent visions of agentic SE (e.g., (Hassan et al., 2025; Roychoudhury et al., 2025)), this paper proposes a ‘whole of processvision that accounts for human and agent actors, artefacts, and activities and other socio-technical concerns across all areas of agentic SE: traditional SE ones renewed for agents and multi-agent systems and new ones for human-agent ecosystems. It presents a preliminary set of guiding values and principles for agentic SE to be comprehensive, responsible, adaptive, foundational, and translational. It offers guidance on designing and using agentic SE vocabulary and welcomes community collaboration to progress these ideas further. Inevitably, proposed ideas end up being defined by practice (e.g., Waterfall’s sequential fate). We may not know what agentic SE truly looks like until it is studied empirically ‘in the wild’. Longitudinal studies to identify useful patterns in agentic repositories and forums and in-depth studies based in interviews, surveys, and experiments with agentic SE teams are needed.
Acknowledgment The author is grateful to Prof John Grundy and Hashini Gunatilake for their valuable feedback and inputs.

### v2

@@
 As technologists rush head-along to make agentic AI a reality, SE researchers are driven to establish agentic SE as a research area.
-While early visions of agentic SE are primarily focused on code-related activities, early empirical evidence calls for a consideration of a range of socio-technical concerns to make it work in practice.
-This paper contributes to the emerging community vision by: (a) recommending an expansion of its scope beyond code, toward a ‘whole of process’ vision, grounding it in SE foundations and evolution and emerging agentic SE frameworks, (b) proposing a preliminary set of values and principles to guide efforts, and (c) sharing guidance on designing/using well-defined vocabulary for agentic SE.
-It is hoped that these ideas will encourage community collaborations and steer the SE community towards laying strong foundations of agentic SE so its not only inevitable but also deliberate and desirable in the long run.
+While early visions of agentic SE are primarily focused on code-related activities, early empirical evidence calls for a consideration of a wider range of socio-technical activities and concerns to make it work in practice.
+This paper contributes to the emerging visions by: (a) recommending an expansion of its scope beyond code, toward a ‘whole of process’ vision, grounding it in SE foundations and evolution and emerging agentic SE frameworks, (b) proposing a preliminary set of values and principles to guide community efforts, and (c) sharing guidance on designing and using well-defined vocabulary for agentic SE.
+It is hoped that these ideas will encourage collaborations and steer the SE community toward laying strong foundations of agentic SE so it is not limited to enabling coding acceleration but becomes the next process-level paradigm shift.
 Agentic software engineering, process, vision, values, principles, vocabulary, terminology, Agentic AI
-This version is the authors’ preprint, submitted to arXiv on 22-Oct-2025.
 1. Introduction
-The Software Engineering (SE) industry, and consequently the SE research discipline, is at the crux of an incoming seismic paradigm shift proposed by the highly anticipated advent of agentic AI.
-Latest releases of agentic systems (e.g., SWE-agent, Passerine, MASAI AutoDev, Devin) are fast making agentic AI a reality.
+The Software Engineering (SE) industry, and consequently the SE research discipline, is at the verge of an incoming seismic paradigm shift proposed by the highly anticipated advent of agentic AI.
+Latest releases of agentic systems (e.g., SWE-agent, Jules, Codex, Devin) are fast making agentic AI a reality.
 Following suit, visions capturing various components of agentic SE are being proposed, e.g., agentic AI software engineers (Roychoudhury et al., 2025), unified software engineering agent (USEagent) (Applis et al., 2025), and structured agentic software engineering (SASE) (Hassan et al., 2025).
@@
 have made AIDev dataset comprising of over 450K agentic pull requests created by leading autonomous coding agents publicly available (Li et al., 2025), positioning agentic SE as an “unfolding reality”.
-These visions, while necessary and welcome, are primarily focused on one SE activity – programming.
-If we expand our vision to focus on SE activities beyond the necessary technical foundations, a range of equally critical socio-technical concerns emerge.
-Early empirical studies show AI’s role as a “personal accelerator’ for coding, writing, and documentation tasks but highlight its limitations in addressing teamwork, coordination, accountability, and culture (Xiao and et al., 2025).
-Similarly, cross-phase applicability, organizational adoption challenges, and human-AI collaboration have been highlighted as areas deserving research focus (Akbar et al., 2025).
+These visions, while necessary and welcome, primarily focus on and around one SE activity – coding, e.g., code generation, review, debugging, repair, configuration etc.
+Whereas other parts of the SE lifecycle, e.g., requirements engineering, design, operations, and practical aspects such as adoption, teamwork, management, and workflow integration also need attention.
+Accepted to the AGENT workshop at the International Conference on Software Engineering (ICSE) 2026
+Early empirical studies show AI’s role as a “personal accelerator” for coding, writing, and documentation tasks but highlight its limitations in addressing teamwork, coordination, accountability, and culture (Xiao et al., 2025).
+Similarly, organizational adoption challenges and human-AI collaboration have been highlighted as areas deserving research focus (Akbar et al., 2025).
 In a review of AI agentic programming, Wang et al.
-highlight the increasing human and socio-technical concerns requiring interdisciplinary solutions (Wang et al., 2025).
+highlight the increasing human and socio-technical concerns requiring interdisciplinary solutions (Wang et al., 2025b).
 Abrahão et al.
 pose several socio-technical questions for the age of AI (e.g., role of human software engineers, how human-AI collaboration might work, how “citizen engineers” might interact with AI agents) (Abrahão et al., 2025).
-There is an urgent need to expand our emerging vision of agentic SE beyond code to examine a full range of SE activities, artefacts, roles, and concerns – traditional ones renewed for agents and multi-agent systems and new ones required to make human-agent ecosystems work.
-This paper contributes to the emerging agentic SE vision by: (a) extending its scope beyond coding, toward a ‘whole of process’ vision, by grounding it in a discussion of SE foundations and evolution (in Section 2), (b) proposing a preliminary set of values and principles to guide our collective design of the emerging vision (in Section 4), and (c) highlighting the importance of, and sharing guidance on, building well-defined vocabulary (terminology) for the emerging agentic SE area (in Section 5).
-Together, it is hoped that these ideas will steer our emerging community vision of agentic SE in the direction of laying the core foundation of agentic SE process models, more fleshed out values and principles, and well-defined vocabulary to support our scholarly conversations, in verbal and written forms.
+For agentic SE to become the next process-level paradigm shift (and not just a coding accelerator), we need to expand the emerging vision of agentic SE beyond code to examine a full range of socio-technical activities, artefacts, roles, and concerns across the entire SE lifecyle, with traditional ones renewed for agents and multi-agent systems as well as new ones required to make human-agent ecosystems work.
+This paper contributes to the emerging agentic SE vision by: (a) extending its scope beyond coding, toward a ‘whole of process’ vision, grounding it in a discussion of SE foundations and evolution (in Section 2 and 3), (b) proposing a preliminary set of values and principles to guide our collective design of the emerging vision (in Section 4), and (c) highlighting the importance of, and sharing guidance on, building well-defined vocabulary (terminology) for the emerging agentic SE discipline (in Section 5).
+Together, it is hoped that these ideas will steer our emerging community vision of agentic SE toward laying the core foundations of agentic SE process models, more fleshed out values and principles, and well-defined vocabulary to support our conversations, in verbal and written forms.
 As with other recent visions, these are not meant to be exhaustive or final.
-Rather, they are intended to encourage community feedback and discussions about adopting a systematic and principled approach to thinking about, researching, and enabling agentic SE.
+Rather, they are intended to encourage community feedback and discussions about adopting a systematic and principled approach to thinking about, researching, and enabling agentic SE (i.e., agentic AI for SE, SE for agentic AI).
 2. SE Foundations and Evolution
 Before we can redefine a field, it makes sense to revisit its foundational definition and evolution.
-SE is defined as “the application of a systematic, disciplined, quantifiable approach to the development, operation, and maintenance of software; that is, the application of engineering to software.” (ISO/IEC/IEEE 24765:2017 Systems and Software Engineering — Vocabulary/Guide to SWEBOK 4.0a [^1] https://ieeecs-media.computer.org/media/education/swebok/swebok-v4.pdf)
-Ergo, SE is more than programming.
+SE is defined as “the application of a systematic, disciplined, quantifiable approach to the development, operation, and maintenance of software; that is, the application of engineering to software.” (ISO/IEC/IEEE 24765:2017 Systems and Software Engineering — Vocabulary/Guide to SWEBOK 4.0a.
+Ergo, SE is more than coding.
 Starting out as a “cottage industry” in the 1950s, SE became an established engineering discipline with many SE process models (AKA software development lifecycle (SDLC), frameworks, methods) proposed, adopted, established, and replaced over the decades (Hoda et al., 2018).
 All of them featured a ‘whole of process’ approach.
-Royce introduced structure and formality to SE in 1970 through a two iteration approach called Waterfall.
+Royce introduced structure and formality to SE in 1970 through a two-iteration approach called Waterfall.
 It was widely adopted as a specification-driven, sequential approach in practice, featuring core SE activities as phases (requirements, analysis, design, coding, testing, and operations), artefacts (documents from each phase), and formalized roles (requirements engineers, designers, developers, testers etc.).
@@
 While, the Rational Unified Process (RUP) highlighted software project management concerns such as timely delivery and budget (Kruchten, 2004).
-Agile software development marked a paradigm shift in SE with its people-centeredness, collaborative approach, focus on working software, and ability to pivot in response to changes (Manifesto, 2001).
+Agile software development marked a paradigm shift in SE with its people-centeredness, collaborative approach, focus on working software, and ability to pivot in response to changes.
 Agile methods, eXtreme Programming and Scrum, came to dominate the SE industry (Hoda et al., 2018).
@@
 Extending the coverage of the SDLC, DevOps connected development and operations.
-Agile became a default/standard (Digital.ai, 2024).
-With the recent resurgence of AI, there was renewed interest in augmenting SE with AI support.
-Dam et al.
-(Dam et al., 2019), proposed a conceptual framework of AI/ML powered agile.
-Building on that, Hoda et al.
-proposed Augmented Agile, an AI-powered agile project management approach, combining the ‘human heart’ (e.g., humans values and concerns) with ‘AI brains’, imagined as an agile co-pilot acting as a team member (Hoda et al., 2023).
-However, these efforts remain conceptual, without application in industrial contexts.
-They foreshadowed what became GenAI/LLM powered SE practices, such as LLM-based multi-agent systems for SE (He et al., 2025).
-In a review of 395 papers focused on the use of LLM-based solutions for SE challenges, Hou et al.
-found a strong focus on individual SE activities, with software development, (programming or coding), being the most researched SE activity (Hou et al., 2024).
-Recently, Zhang et al.
-proposed AgileGen, an agile-based generative software development through human-AI teamwork framework, focused on maintaining consistency between the generated software and user requirements (Zhang and et al., 2025).
-While others have called for GenAI to be human-centered (Russo et al., 2024).
+Agile became an industry default/standard (State of Agile report, 2025).
+The use of intelligent techniques to support various activities of the SE lifecycle has been of interest over the years (Perkusich et al., 2020).
+In 2019, a conceptual framework of AI/ML powered agile was proposed (Dam et al., 2019).
+Building on that, a vision of Augmented Agile, an AI-powered agile project management approach, combining the ‘human heart’ (e.g., humans values and concerns) with ‘AI brains’, imagined as an agile co-pilot team member was proposed (Hoda et al., 2023).
+These techniques and conceptual ideas foreshadowed GenAI/LLM powered SE practices, such as LLM-based multi-agent systems for SE (He et al., 2025).
+A review of 395 papers on the use of LLM-based solutions for SE found that software development (i.e., coding and related) was the most researched activity while human-intensive activities (e.g., RE, design, management) were the least explored (Hou et al., 2024).
+Recently, AgileGen, focused on maintaining consistency between the generated software and end-user requirements was proposed (Zhang et al., 2025).
+Others have called for GenAI to be human-centered (Russo et al., 2024).
 Figure 1. Software Engineering Evolution: A Brief Timeline
-There is an urgent need to expand our vision of agentic SE beyond code (design, writing, testing, reviewing, repairing) to examine a full range of agentic SE areas and associated actors, activities, artefacts, and other socio-technical concerns – traditional SE ones renewed for agents and multi-agent systems and new ones needed to make human-agent ecosystems work.
+Software Engineering Evolution: A Brief Timeline
+For agentic SE to become the next process-level paradigm shift in SE, we need to expand our vision of agentic SE beyond code to examine a full range of areas and activities (e.g., ethical alignment, RE, design, development, operations), associated actors (e.g., teams of human and AI software engineers, end-users), artefacts (e.g., code, data, evidence and analyses samples, experimental infrastructure), and other socio-technical concerns – traditional SE ones renewed for agents and multi-agent systems, and new ones needed to make human-agent ecosystems work.
 The establishment of SE can be credited to the formalization of various process models, the development of supporting tools, techniques, and the formalization of knowledge areas and vocabulary.
@@
 3. Toward Agentic SE (emerging)
-Agentic AI draws its foundations from agent-oriented software engineering (Wooldridgey and Ciancarini, 2000; Hassan et al., 2025; Roychoudhury, 2025).
-Introduction of agentic systems such as SWE-agent, Google’s Passerine, Microsoft’s MASAI, AutoDev, Devin are making agentic AI a reality.
+Google’s Jules, OpenAI’s Codex, Cognition’s Devin, SWE-agent, AutoCodeRover, and Anthropic’s Claude Code are making agentic AI a reality.
 Unsurprisingly, visions of new agentic SE frameworks and approaches are being proposed.
@@
 (Li et al., 2025).
-These visions and efforts, while necessary and welcome, are primarily focused on one SE activity – programming.
+These visions and efforts, while necessary and welcome, are primarily focused around one SE activity – coding.
 Besides the technical challenges of implementing agentic AI for programming-related SE tasks, a number of socio-technical concerns across the SDLC arise.
-Early empirical evidence highlight the need to address agentic AI applicability across the SE process.
-Xiao et al.’s longitudinal study showed AI’s role as ‘a personal accelerator’ but it was not seen to fix teamwork issues and its impact on coordination, accountability, and culture remain to be seen (Xiao and et al., 2025).
+Early empirical evidence highlights the need to address agentic AI applicability across the SE process.
+Xiao et al.’s longitudinal study showed AI’s role as ‘a personal accelerator’ but it was not seen to fix teamwork issues and its impact on coordination, accountability, and culture remain to be seen (Xiao et al., 2025).
 Akbar et al.
@@
 Similarly, Wang et al.
-review AI agentic programming and highlight that increasing human and socio-technical concerns with AI agentic programming such as context handling, human-AI collaboration, verification of agent behaviour, will require “interdisciplinary solutions(Wang et al., 2025).
+review AI agentic programming and highlight that increasing human and socio-technical concerns with AI agentic programming such as context handling, human-AI collaboration, verification of agent behaviour, will require “interdisciplinary solutions(Wang et al., 2025b).
 Similar socio-technical issues (e.g., human-AI collaboration, next generation education, job security, ethics, diversity) were highlighted by Abrahão et al.
 (Abrahão et al., 2025).
-How nascent agentic SE is can be gauged from the fact that most of these proposals are under review (Oct 2025).
+How nascent agentic SE is can be gauged from the fact that most of these proposals are under review (Jan 2026).
 A ‘Whole of Process’ Vision
-Building on the foundational aspects of previous SE models (Section 2) and AI capabilities captured in proposed frameworks above, this paper presents a preliminary ‘whole of process’ vision for agentic SE (Figure 2).
+Agile was a process revolution, defining new SE roles, practices, and artefacts (Hoda et al., 2018).
+Agentic SE is a technology revolution.
+Currently, there is no guidance on how AI teammates can/should be incorporated across the SDLC.
+Building on the foundational aspects of previous SE models (Section 2) and AI capabilities of the proposed frameworks (Section 3), this paper presents a preliminary ‘whole of process’ vision for agentic SE (Figure 2).
 It is not meant to be a definitive agentic SE process model, rather a ‘first look’ that will need to be improved and expanded on.
-The vision extends across high-level core areas, approached iteratively, by human and agent (single AI agents or a multi-agent system) actors, collaborating in activities at varying levels of AI agency (depicted by the slider under human control).
-  • Agentic SE Ethical Alignment: covers activities, actors, and artefacts (AAA) to do with AI and software ethics alignment.
-With the increasing capabilities and agency of AI models and their potential for wide-scale impact, ethical alignment is a must.
-  • Agentic Requirements Engineering: covers AAA to do with software requirements (customer, user, subject matter expert, quality, and ethical requirements)
-  • Agentic Design: covers AAA to do with software design (architecture, UI, UX).
-  • Agentic Development: covers AAA to do with software code design, production, testing, reviewing, repairing.
-We have a head-start on envisioning the transformation in this area with emerging visions and an agentic dataset.
-(Roychoudhury et al., 2025; Roychoudhury, 2025; Hassan et al., 2025; Li et al., 2025).
+The vision extends across high-level core areas of the SE lifecycle, approached iteratively, by human and agent (single agents or multi-agent system) actors, collaborating in activities at varying levels of AI agency (depicted by the slider under human control (Spiegler et al., 2025)).
+Given AI’s increasing impact on society, it introduces ethical alignment as first and central.
+  • Agentic SE Ethical Alignment: covers activities, actors, and artefacts (AAA) to do with ethics alignment.
+With increasing AI capabilities and agency and their potential for wide-scale societal impact (Spiegler et al., 2025), alignment with safe AI laws and regulations at the micro (e.g., project, company) and macro (e.g., domain, region, country, international) levels is a must.
+  • Agentic Requirements Engineering: covers AAA to do with software requirements (e.g., customer, user, subject matter expert, quality, and cascading ethical requirements.)
+  • Agentic Design: covers AAA to do with software design (e.g., architecture, UI, UX, design for human benefit and equity).
+  • Agentic Development: covers AAA to do with software code design, production, testing, reviewing, repairing, etc.
+We have a head-start on envisioning the transformation in this area with emerging visions and an agentic AI dataset (Roychoudhury et al., 2025; Roychoudhury, 2025; Hassan et al., 2025; Li et al., 2025).
   • Agentic Operations: covers AAA to do with deployment, maintenance, and operations (e.g., Ops/DevOps/DevSecOps).
-These high-level descriptions above need inputs from relevant area experts to develop details of how the vision may be implemented in practice.
-Transforming relevant SE areas and introducing new ones (e.g., ethical alignment) will ensure agentic SE leverages the full capabilities of agents to support humans and improve outcomes across the process, not only to do with code-related activities.
+These high-level area descriptions will need to be fleshed out in collaboration with area experts, including relevant RQs, example scenarios, and specific challenges.
+Transforming traditional SE areas and introducing new ones (e.g., ethical alignment) will ensure agentic SE leverages the full capabilities of agents to support humans and improve outcomes across a full range of SE activities.
 Figure 2. A ‘whole of process’ vision for Agentic SE, where human and agent actors iterate across areas, activities, artefacts with varying levels of AI agency (under human control)
+A ‘whole of process’ vision for Agentic SE
 4. Toward Agentic SE Values and Principles
-To guide this paradigm shift, a set of preliminary CRAFT values and principles for agentic SE is presented to spark community conversation and collaboration.
+To guide this paradigm shift, a set of preliminary comprehensive, responsible, adaptive, foundational, and translational (CRAFT) values and principles for agentic SE is presented to spark community conversation and collaboration.
+Most apply to both research ( ) and practice ( ), while some are primarily research focused.
 Value 1. Comprehensive:
@@
 Comprehensive Whole of Process Approach Agentic SE needs to take a ‘whole of process’ approach to examining traditional SE areas and associated activities, actors, artefacts (AAA) and consider new ones to enable emerging human-agent ecosystems.
-Human and Socio-technical Aspects Agentic SE needs to include a focus on human and socio-technical aspects of SE including roles played by human and agent actors, human aspects that apply to humans and agents (e.g., human values, empathy, motivation, emotional intelligence, personality) and socio-technical aspects (e.g., trust, adoption barriers, teamwork, collaboration, communication, coordination).
-Responsible Ethics-by-Design Agentic SE needs to needs to apply an ethics-by-design approach to its design and application, questioning the why, and not being driven by the why not.
-Sustainability Agentic SE need to ensure sustainability of the core technologies, ecosystems, and the planet.
-Adaptive Industrial Relevance Agentic SE need to adapt to new AI models and paradigms and work closely with the AI community and co-design with the industry.
+Human, Agentic, and Socio-technical Aspects Agentic SE needs to include a focus on human, agentic, and socio-technical aspects of SE including roles played by human and agent actors, human and agentic aspects (e.g., human values, empathy, motivation, emotional intelligence, personality) and socio-technical aspects (e.g., mutual trust, adoption barriers, teamwork, collaboration, communication, coordination, project management).
+Responsible Ethics-by-Design Agentic SE needs to apply an ethics-by-design approach to its design and application, questioning the why, and not being driven by the why not.
+Sustainability Agentic SE needs to ensure sustainability of the core technologies, ecosystems, and the planet.
+Adaptive Industrial Relevance Agentic SE needs to adapt to new AI models and paradigms and work closely with the AI community and co-design with the industry.
 Socio-technical Relevance Agentic SE needs to predict and report socio-technical impacts of changes in AI models and paradigms.
-Foundational Foundational Knowledge Agentic SE need to develop foundational knowledge about AAA across all areas, not only programming.
-Foundational Solutions Agentic SE need to develop foundational solutions to support AAA across all areas, not only programming.
-Translational Awareness and Education Agentic SE need to build awareness through education, certifications, funding bodies, media, and society; and maintain incident logs.
-Actionable Guidelines Agentic SE need to study and enable the adoption of research knowledge and solutions into practice through usable guidelines and tools.
+Foundational Foundational Knowledge Agentic SE needs to develop foundational knowledge about AAA across all areas, not only coding/programming.
+Foundational Solutions Agentic SE needs to develop foundational solutions to support AAA across all areas, not only coding/programming.
+Translational Awareness and Education Agentic SE needs to build awareness through education, certifications, funding bodies, media, and society; and maintain incident logs.
+Actionable Guidelines Agentic SE needs to study and enable the adoption of research knowledge and solutions into practice through usable guidelines and tools.
   • Principle of Whole of Process Agentic SE needs to take a ‘whole of process’ approach – A ‘whole of process’ approach, a hallmark of previous SE models (section 2).
@@
 Relevant parts of the SE lifecycle, not only its code-related activities, need to be considered as well as new required activities, in designing agentic SE.
-This includes assessing the goals and constrains related to ideation, RE activities (e.g., to establish user needs and preferences), design, architecture, project management (e.g., costs, times, resources), functional and user testing, delivery, maintenance, operations/evolution, as well as to new practices, trade-offs, and process optimization.
-  • Principle of Human and Socio-technical AspectsAgentic SE needs to include the human and socio-technical aspects of SE – Like SE, agentic SE is socio-technical, where “the social and technical aspects are inherently interwoven(Hoda, 2024).
+This includes assessing the goals and constrains related to ideation, RE activities (e.g., to establish user needs and preferences), design, architecture, project management (e.g., costs, times, resources), functional and user testing, delivery, maintenance, operations/evolution, as well as for new practices, trade-offs, and optimizations.
+  • Principle of Human, Agentic, and Socio-technical Aspects Agentic SE needs to include the human, agentic, and socio-technical aspects of SE – Like SE, agentic SE is socio-technical, where “the social and technical aspects are inherently interwoven(Hoda, 2024).
 Technical innovation is necessary but not sufficient for advancing agentic SE, e.g., trust and utility concerns outrank performance in real-world adoption (Li et al., 2025).
-Human and socio-technical aspects must be considered, e.g., the roles played by human and agent actors, human aspects (e.g., human values, empathy, emotional intelligence), and socio-technical aspects (e.g., trust, adoption barriers, human experience, coordination, communication, collaboration, conflict management, teamwork).
+Human, agentic, and socio-technical aspects must be considered, e.g., the roles played by human and agent actors, human and agentic aspects (e.g., human values, empathy, emotional intelligence), and socio-technical aspects (e.g., mutual trust, adoption barriers, human experience, coordination, communication, collaboration, conflict management, teamwork).
 Research on these will help the community answer fundamental questions such as: What should be automated, to what extent (human oversight levels), and how?
@@
 Value 2. Responsible:
-Agentic SE needs to ensure ethics-by-design and sustainability. With agentic AI injecting more agency into AI, we need to aim for positive impact and human control of AI (Spiegler et al., 2025).
+Agentic SE needs to ensure ethics-by-design and sustainability. With agentic AI injecting more agency into AI, we need positive societal impact and human control of AI (Spiegler et al., 2025).
   • Principle of Ethics-by-Design Agentic SE needs to apply an ethics-by-design approach to its design and application – There is unprecedented growth in AI-based companies (e.g., OpenAI, NVIDIA, Microsoft, Cursor) with claims of 100x productivity [^1] https://www.businessinsider.com/surge-ceo-ai-100x-engineers-2025-7.
@@
 why not automate X).
-  • Principle of SustainabilityAgentic SE needs to ensure sustainability – The planetary cost of AI technologies is non-trivial (Crownhart, 2025).
+  • Principle of Sustainability Agentic SE needs to ensure sustainability – In addition to ensuring ultimate human control and positive social impact (Spiegler et al., 2025), the planetary cost of AI technologies cannot be ignored (Crownhart, 2025).
 As Cruz et al.
-highlight in their article on Green AI, “our [SE] community needs to adopt a long-term plan enabling a conscious transformation that aligns with environmental sustainability values(Cruz et al., 2025), which requires handling trade-offs (Lago et al., 2015).
+highlight in their article on Green AI, “our [SE] community needs to adopt a long-term plan enabling a conscious transformation that aligns with environmental sustainability values(Cruz et al., 2025), which necessitates handling trade-offs (Lago et al., 2015).
 Value 3. Adaptive:
 Agentic SE needs to adapt to maintain industrial and socio-technical relevance.
-  • Principle of Industrial RelevanceAgentic SE researchers need to adapt to new AI models and paradigmsAgentic SE fundamentally ties SE to AI.
+  • Principle of Industrial Relevance Agentic SE researchers and practitioners need to adapt to new AI models and paradigmsAgentic SE fundamentally ties SE to AI.
 Agentic SE and AI communities should work in tandem, adapting and contributing to the latest AI models and paradigms.
 Empirical studies of problem and solution spaces will be useful in maintaining industrial relevance.
-Co-designing and evaluations with industry will also help.
+Co-designing and evaluations with industry will help.
   • Principle of Socio-technical Relevance Agentic SE needs to predict and report socio-technical impacts of changes in AI models and paradigms – Maintaining relevance extends to the associated human (e.g., evolving levels of AI agency and human control (Spiegler et al., 2025)) and socio-technical aspects (e.g., awareness, ethics, adoption).
@@
 Value 4. Foundational:
-Agentic SE needs to develop foundational knowledge and solutions.
-  • Principle of Foundational KnowledgeAgentic SE needs to develop foundational knowledge – Like how agile established foundational knowledge through empirical studies of new roles, practices, artefacts, and socio-technical aspects (Hoda et al., 2018), a similar approach will benefit agentic SE.
-Building original taxonomies and theories (Hoda, 2024) and testing existing ones will contribute to theoretical foundations, e.g., agents in SE taxonomy (Wang and et al., 2025).
-Qualitative (Seaman et al., 2025), quantitative, and mixed methods studies (Storey et al., 2025) will help build a body of knowledge.
-  • Principle of Foundational SolutionsAgentic SE needs to develop foundational solutions – With proposed visions of programming frameworks (Hassan et al., 2025; Roychoudhury et al., 2025), there is already movement on this front.
-As with GenAI/LLM based solutions (technical tools, techniques, frameworks, platforms) for SE (Hou et al., 2024), we can expect foundational solutions for agentic SE being proposed by the SE community.
-It will be critical to extend the solutions to support other parts of agentic SE areas beyond programming.
+Agentic SE needs to develop foundational knowledge and solutions beyond coding.
+  • Principle of Foundational Knowledge Agentic SE needs to develop foundational knowledge beyond coding – As with agile methods (Hoda et al., 2018), establishing foundational knowledge through empirical studies of new roles, practices, artefacts, and socio-technical aspects will benefit agentic SE.
+Building original taxonomies and theories (Hoda, 2024) and testing existing ones will contribute to theoretical foundations, e.g., agents in SE taxonomy (Wang et al., 2025a).
+Qualitative (Seaman et al., 2025), quantitative, and mixed methods studies (Storey et al., 2025), and intra/inter-disciplinary work will help build a body of knowledge.
+  • Principle of Foundational Solutions Agentic SE needs to develop foundational solutions beyond coding – Proposed visions of programming frameworks are emerging (Hassan et al., 2025; Roychoudhury et al., 2025).
+As with GenAI/LLM based solutions (tools, techniques, frameworks, platforms) for SE (Hou et al., 2024), foundational solutions for agentic SE are expected to be proposed.
+It will be critical to extend such solutions to support other agentic SE areas beyond coding/programming.
 Value 5. Translational:
 Agentic SE needs to enable research translation into practice through building awareness, education, and actionable guidelines.
-  • Principle of Awareness and EducationAgentic SE researchers need to build awareness – While awareness of agentic AI is on the rise thanks to agentic AI systems, products, and news (including ethical incidents), awareness of agentic SE will need to be built through education and training within and across disciplines (in classrooms and workplaces) as well as with accreditation, certification, funding bodies, media, and society.
+  • Principle of Awareness and Education Agentic SE needs to build awareness – While awareness of agentic AI is on the rise thanks to agentic AI systems, products, and news (including ethical incidents), awareness of agentic SE (as a discipline) will need to be built through education and training (in classrooms and workplaces) as well as with accreditation, certification, funding bodies, media, and society.
 Maintaining an agentic SE incident log will also help.
-Previous SE models have shown the importance of such efforts to the long-term future of those models and their associated communities.
-  • Principle of Usable Guidelines and ToolsAgentic SE researchers need to study and enable the adoption of research knowledge and solutions into practice – Building knowledge and solutions is necessary but not sufficient to progress agentic SE.
-To achieve real-world impact, they need to be translated into usable guidelines and tools for practice.
-Mirroring and leveraging AI’s multi-modal generation, they can have more industry-friendly interfaces and mode of delivery (e.g., posters, executive summaries, graphical abstracts, short videos).
+Previous SE models have shown the importance of such efforts to the long-term future of those process models and their associated communities.
+  • Principle of Usable Guidelines and Tools Agentic SE needs to study and enable the adoption of research knowledge and solutions into practice – Building knowledge and solutions is necessary but not sufficient to progress agentic SE.
+To achieve real-world impact, they need to be translated into actionable guidelines and tools for practice.
+Mirroring and leveraging AI’s multi-modal generation, they can have more industry-friendly interfaces and modes of delivery (e.g., posters, executive summaries, graphical abstracts, short videos).
 5. Toward Agentic SE Vocabulary
@@
 distinguish between AI agents and agentic AI and present a taxonomy of AI agent paradigms (Sapkota et al., 2025)) and ultimately, knowledge areas.
-While agentic SE is nascent, discrepancies in terminology usage are already apparent.
+Although agentic SE is nascent, drifts and discrepancies in terminology usage are already apparent.
 Some are syntactic (e.g., agentic AI software engineer vs.
@@
   • Relevancefit for purpose and domain context – to what extent is the term relevant to the domain? E.g., traditional SE vocabulary can be revisited for relevance in the agentic SE context.
-  • Coverageconceptual completeness and accuracy – to what extent does the term cover (nearly) all the goals and responsibilities of the agentic SE process area, actor, activity, or artefact it aims to support or automate?
-For example, a software engineer is more than a programmer, so can an agent that enacts programming tasks be termed a software engineer?
-Autonomous coding agent (Li et al., 2025) is a more accurate description.
+  • Coverageconceptual completeness and accuracy – to what extent does the term cover (nearly) all the goals and responsibilities of the agentic SE process area, actor, activity, or artefact it aims to support?
+For example, a software engineer is more than a coder, so an autonomous coding agent (Li et al., 2025) needs to be able to do more than coding to be termed an AI software engineer.
   • Acceptancerecognition and adoption by the community – to what extent will humans (especially in the role) agree to an AI agent being given their professional title?
 For example, people might judge human and robot paintings as art but are less willing to see robots as artists (Mikalonytė and Kneer, 2022).
-  • Consistencyuniformity of use, stability, adaptability – to what extent do different research groups around the world within SE use the terminology in the same way, both in form (syntax) and meaning (semantics)?
+  • Consistencyuniformity of use, stability, adaptability – to what extent do different technology and research groups in the SE world use the terminology in the same way, both in form (syntax) and meaning (semantics)?
 how long do they remain stable?
-when and how do they adapt to changes over time?
+how do they adapt to changes over time?
   • Philosophical alignmentreflexivity leading to awareness of underlying assumptions, power dynamics, fairness – reflecting on who comes up with terminology and why?
@@
 Building on the foundational aspects of previous SE models (Section 2) and recent visions of agentic SE (e.g., (Hassan et al., 2025; Roychoudhury et al., 2025)), this paper proposes a ‘whole of processvision that accounts for human and agent actors, artefacts, and activities and other socio-technical concerns across all areas of agentic SE: traditional SE ones renewed for agents and multi-agent systems and new ones for human-agent ecosystems.
-It presents a preliminary set of guiding values and principles for agentic SE to be comprehensive, responsible, adaptive, foundational, and translational.
-It offers guidance on designing and using agentic SE vocabulary and welcomes community collaboration to progress these ideas further.
-Inevitably, proposed ideas end up being defined by practice (e.g., Waterfall’s sequential fate).
+It presents a preliminary set of comprehensive, responsible, adaptive, foundational, and translational (CRAFT) values and principles for agentic SE.
+It offers guidance on designing and using agentic SE vocabulary, welcoming community collaboration to progress ideas.
+Inevitably, proposed ideas end-up being defined by practice (e.g., Waterfall’s sequential fate).
 We may not know what agentic SE truly looks like until it is studied empirically ‘in the wild’.
-Longitudinal studies to identify useful patterns in agentic repositories and forums and in-depth studies based in interviews, surveys, and experiments with agentic SE teams are needed.
-Acknowledgment The author is grateful to Prof John Grundy and Hashini Gunatilake for their valuable feedback and inputs.
+Longitudinal studies to find useful patterns in agentic repositories and forums, and in-depth empirical studies based on interviews, surveys, and experiments with agentic SE teams, using inter- and intra-disciplinary approaches, are needed.
+Acknowledgment Thanks to John Grundy, Hashini Gunatilake, Abhik Roychoudhury, and the reviewers for their valuable feedback.